OpenStudy (anonymous):

(a+b)^2y^2-4aby-(a-b)^2 factorize.??

OpenStudy (freckles):

\[(a+b)^2 \cdot y^2 -4ab \cdot y -(a-b)^2 \\ A=(a+b)^2 \\ B=-4ab \\ C=-(a-b)^2 \\ \text{ where } Ay^2+By+C \\ A \cdot C =-(a+b)^2(a-b)^2 \text{ and again } B=-4ab \\ \text{ We want to find two factors of } -(a+b)^2(a-b)^2 \\ \text{ that have product } -(a+b)^2(a-b)^2 \\ \text{ and sum } -4ab \]

OpenStudy (freckles):

if that is possible..

OpenStudy (freckles):

We could probably cheat and used the quadratic formula and then go backwards from there.

OpenStudy (freckles):

But I want to see if we can think of the first way first

OpenStudy (anonymous):

could you plz simplify it for me..!!

OpenStudy (freckles):

Oh I thought you wanted to factor it?

OpenStudy (freckles):

I don't if you want to factor or if you want to multiply it out (since you said simplify)

OpenStudy (freckles):

Factor is to write it as something*something

OpenStudy (freckles):

Can you please define simplify @abaham ?